Sol Summary – January 25th

Sol: 4
Summary Title: Expanding the Search
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Sol 4 had the potential of denying us EVAs based on the weather forecast. Fortunately, the rain held off and we were able to get away on time. Clare had us at Kissing Camel Ridge West in search of concretions (a source of carbonates) and petrified wood (an analogue of organic remains). Previous crews had reported this location and, after some climbing around outcrops and mudstone domes, we located the stratigraphic layer that contains the concretions. GPS coordinates were taken and samples collected. The petrified wood remained elusive as only fragments in the stream wash were observed. The second EVA saw an excursion towards the north along Cow Dung Road past Cowboy Corner en route to the largest stream draining the Tharsis Montes. Operational constraints (Point of No Return is rover power levels at 55%) pulled the team up short of the proposed destination. Nevertheless, Steve was able to have the push broom spectrometer up and recording the many and varied lithologies of the nearby geology. Finally, the commander joined a livestream event for Igniting The Space Down Under timed for transmission for Australia Day 26th January back on Earth.
Look Ahead Plan: Weather dependent as always, tomorrow will see an EVA return to the Tharsis Montes area for the next stage in collecting spectroscopic data. The remaining time will be spent continuing the troubleshooting of the stubborn sensor station and collating previously collected data.
Anomalies in work: Nothing to report.
Weather: Overcast morning and remained so during the day. No precipitation. Temperature increasing from 0 Celsius in the morning to high single digits in the afternoon. Humidity fluctuated between 80% and 50%. Faint to no wind.
Crew Physical Status: nominal
EVA: EVA 7 with Clare, Andrew and Scott to Kissing Camel Ridge West. EVA 8 with Steve, Alex and Rob to the Tharsis Montes area north of Cowboy Corner.
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, EVA Request(s). No Astronomer report (too cloudy for telescopic observations).
Support Requested: Nothing to request.

Copyright © The Mars Society. All rights reserved. | Main Site